OpenStudy (anonymous):

How do I graph \[3(x+3)^2+3\] without using a graphing calculator?

OpenStudy (anonymous):

\[3(x+3)^2+3 \implies 3 \times (x + 3)(x+3)+3\] Use FOIL: \[3 \times (x^2+6x+9)+3\] Multiply using distributive property(simplify): \[3x^2+18x+30\] Now how do I graph that? \[y = mx + c~?\]

OpenStudy (campbell_st):

use a table of values... and a little knowledge its a concave up parabola with a vertex at (-3, 3) so the parabola is postive definite... meaning no zeros hope this helps

OpenStudy (anonymous):

It did, thank you very much sir :-).